OpenStudy (anonymous):

pleaseeee! x^2 + 3x - 28 = 0

OpenStudy (anonymous):

I just gave you a site to help you with that. The site had step by step instructions. Stop being lazy we will not do your hw for you. We are here to help.

OpenStudy (anonymous):

i was confused.

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

at the x^2 + 3x - 28 = 0 there is the three x and according to my instructions i should move the 28 to the outside. and then divide the three in half and square root that. then having x^2 + 3x + 1/4 = 28 + 1/4

OpenStudy (anonymous):

okay. wel i hope someone else helps me then since ur being rude.